Determine the pH of a solution created by adding 350.0 mL of 1.25 mol/L HBr (aq) to 750.0 mL of 1.00 mol/L Sr(OH)2 (aq)


1
Expert's answer
2021-03-22T06:28:05-0400

HBR + Sr(OH)2 > SrBr^2 +H20

New volume = (350+750)= 1000ml

n(Sr(OH)2= Mv = (1×750)(10^-3L/ml) = 7.5 ×10^-1mol

n(HBr) = Mv = (1.25×350)(10^-3L/ml) = 4.375×10^-1 mol

New Sr(OH)2 = n/Vf= (3.125×10^-1mol)(1000ml)(10^-3L/ml) = 0.3125

New SrBr^2 = n/vf = (4.374×10^-1mol)(1000ml)(10-3)

= 0.4375

= -log(0.3125/0.7142)

= -0.1461